OpenStudy (anonymous):

Factor and Simplify: 4x^2-64 3x -------- x ------ x^3 8x+32

OpenStudy (lgbasallote):

where are you stuck there?

OpenStudy (anonymous):

The whole thing... /:

OpenStudy (lgbasallote):

you don't know how to start?

OpenStudy (anonymous):

Nope

OpenStudy (lgbasallote):

hmm do you know how to factor expressions?

OpenStudy (anonymous):

Barely

OpenStudy (lgbasallote):

do you know how to factor out 4x^2 - 64?

OpenStudy (anonymous):

would it be 4x^2=400-64? I feel like that's completely wrong

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (anonymous):

the 400? 4x^2 = 400? 400 - 64 = 336

OpenStudy (anonymous):

is x=10 given?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

Can someone please just help find the answer?

OpenStudy (anonymous):

First you factor out the numbers that you can So in the numerator on the left, you would factor out 4. Then you do the same for the bottom denominator on the right since you can also factor it, so you would factor out 8. Then you could see that (x^2-16) is a special function where you could further factor it into (x+4)(x-4) Then you could do the elimination process of getting rid of the x+4 , and also the number 4 and 8. You can also get rid of x, but only 1 since that is common between them. 4x^2-64/x^3 x 3x/8x+32 = 4(x^2-16)/x^3 x 3x/8(x+4) = 4(x+4)(x-4)/x^3 x 3x/8(x+4) = (x-4)/x^3 x 3x/2 = 3(x-4)/ 2x^2 That is it. I don't think you can simplify it anymore. Hope it helped!

OpenStudy (anonymous):

Than you SO much I think I got it a little. But thank you!!

OpenStudy (anonymous):

No problem (: